OpenStudy (anonymous):

A large bin of mixed nuts contains peanuts, cashews, almonds, and walnuts. Sheila scoops out a bowl full of nuts and counts how many of each type she has : 17 peanuts, 9 cashews, 12 almonds, and 10 walnuts. Armando scoops out a bowl of 32 nuts. About how many should he expect to be cashews? A.5 B.6 C.9 D.18

OpenStudy (anonymous):

17 peanuts, 9 cashews, 12 almonds, and 10 walnuts. Is a random selection. 17+9+12+10=48 9 out of 48 is cashews 9/48=3/16 For another sample we draw 32 nuts Expected to get 32*3/16=2*3=6

OpenStudy (dwelsher):

d

OpenStudy (dwelsher):

9/48=~18or 19 therefore what is you close answer.Use process of elimination to solve